OpenStudy (anonymous):

Sarah Meeham blends coffee for Tasti-Delight. She needs to prepare? 190 pounds of blended coffee beans selling for $4.05 per pound. She plans to do this by blending together a high-quality bean costing $5.00 per pound and a cheaper bean at $3.00 per pound. To the nearest pound, find how much high-quality coffee bean and how much cheaper coffee bean she should blend.

OpenStudy (nurali):

Let x = pounds of the 5 dollar a pound coffee Let y = pounds of the 3 dollar a pound coffee

OpenStudy (nurali):

\[\frac{ 5x+3y }{ 190}=4.05\] and \[x+y=190\]

OpenStudy (nurali):

YES. TWO Equations. They are both linear equations. Work with the ratio equation so it is either easier to use for substitution or so it is in Standard Form

OpenStudy (nurali):

\[5x+3y=190*4.05\] \[5x+3y=769.5\] \[5(190-y)+3y=769.5\]

OpenStudy (nurali):

\[950-5y+3y=769.5\] \[-2y=-180.5\] \[y=90\] and \[x=100\]

OpenStudy (nurali):

x=100 pounds, the more expensive coffee y=90 pounds, the cheaper coffee